OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 3x + 6y = 36 3x - 6y = 0

OpenStudy (anonymous):

if you take the equation 3x-6y=0 make it 3x=6y by moving the 6y

OpenStudy (anonymous):

by adding u will get, 6x = 36 x = 36/6 x = 6 3(6)-6y = 0 18-6y = 0 -6y = -18 y = -18/-6 y = 3

OpenStudy (anonymous):

now that you know that 3x and 6y are the same look at the other equation, is theyre the same they both equal 18 3x=18 x=6 6y=18 y=3 so the answer id (6,3)